The width of a rectangle is half its length. If its length is x feet, find its perimeter in terms of x.
Answer by Theo(13342)

length = x
width = x/2
p = 2 * (length + width) = 2 * (x + x/2) = 2x + x = 3x
the perimeter is equal to 3x.
that's your solution.